# Paying to Do Better: Games with Payments between Learning Agents

### v1

@@ 3 Auctions with Payments between Learning Agents
To get a deeper understanding of the incentives of users in online platforms to use payments with their learning agents and the potential impact of such payments, we focus on one class of games that have already been widely analyzed: auctions with automated bidders. We consider the simplest model of first- and second-price auctions, where at each time $t\in[T]$ a single identical item is sold. Each player $i$ has a value $v_{i}$ for the item, and we index the players in decreasing order of their values, $v_{1}\geq v_{2}\geq\dots\geq v_{n}$. At round $t$, each agent submits a bid $b_{i}^{t}$ and the auction mechanism determines the identity of the winner and a price $p^{t}$. The payoff for the winning agent from the auction is $v_{i}-p^{t}$; the other agents get payoff zero. In addition to payoffs from the auction, agents get a payoff according to the payments between the agents (as defined in Section 2). Utilities are additive over game rounds. In the first-price auction, the payment is equal to the highest bid; in the second-price auction, the payment is the second-highest bid. In both cases the highest bidder wins, and we assume tie-breaking according to the index of the agent (as is done, e.g., in [35]), so that if $i<j$ and $i$ and $j$ both have the same top bid, then $i$ is taken to be the winner. All our analyses are similar for arbitrary tie-breaking rules, but breaking ties in favor of the highest-value agent simplifies the presentation. Formal proofs for this section are in Appendix D.
Second-price auctions:
@@
The second-price auction has been widely studied [86] (see Appendix A). While the auction is known to be incentive compatible, it is also known that truthful bidding forms only one out of many Nash equilibria of the stage game. In the context of learning dynamics, it has been shown that regret-minimizing agents in this auction do not converge to the truthful equilibrium; instead, they reach some degree of tacit collusion [53] with lower revenue to the auctioneer than that obtained under truthful bids.
@@
Specifically, simplifying slightly, second-price auctions have two types of non-truthful equilibria. The first type is “overbidding equilibria,” where the high-value player bids anything above his value and the low-value player bids anything below the high value. The second type is “low-revenue equilibria,” where the high-value player bids anything above the low value and the low-value player bids anything below the low value. [^7] Both types of equilibria arise for any $n$; we describe the two-player case for ease of exposition. The latter type is especially interesting for our setting since any mixture of such equilibria is consistent with regret minimization, but yields high welfare to the players. This suggests that when learning agents exchange payments, the high-valued agent may be able to use payments to influence the dynamics with the other agents to induce a better equilibrium for himself.
@@
We show that, indeed, when agents “trade outside of the mechanism” in this way during their dynamics, this kind of collusion emerges from the agents’ interaction in a strong way, potentially leading to zero revenue for the auctioneer.
@@
The single-item second-price auction where the players’ top two values are $v_{1}>v_{2}$ is not stable for any regret-minimizing agents for the players, and the payment-policy game has an $\epsilon$-equilibrium where the players capture the full welfare and the auctioneer gets revenue zero.
@@
Proof sketch: The idea of the proof is, first, to construct a unilateral payment policy for player $1$, parameterized by a single parameter $\epsilon$ that determines the amount that agent $1$ pays to the other agents, such that this policy induces a dynamic where player $1$’s (i.e., the high player’s) agent bids his value and all the other learning agents bid zero. We then show that for every other player, using the zero-payment policy is in fact a best response, and that the (net) payoff for the high-value player can be arbitrarily close to the full welfare, so player $1$ cannot improve his payoff by more than $\epsilon$ for any $\epsilon>0$. (Note, that with the second-price payment rule, if the agents were not restricted to using finite payments, the players would be locked into a game without an equilibrium, where each player tries to commit to making the highest bid and incentivize the others to bid zero.)
First-price auctions:
@@
In contrast to the second-price auction, the first-price auction is known not to be incentive compatible. Intuitively, the direct dependence of the utility on the bid requires players to constantly optimize their responses to the bids of the other players. There has thus been significant interest in learning strategies, outcomes, and dynamics in this setting (see Appendix A for references). Interestingly, despite the significant difference from the truthful second-price auction, it is well known (and not hard to see) that the Nash equilibrium of the first-price auction yields the truthful outcome of the second-price auction. It has been shown that mean-based [19] (see Appendix C) regret-minimization dynamics [53] and best-response dynamics [71] can converge only to this outcome.
@@
In [35], it was shown that the first-price auction also has “collusive” coarse correlated equilibria [^8] An equilibrium notion equivalent to all players having no regret. See [78] and Appendix C. (CCE) with lower revenue for the auctioneer than the second-price outcome, and higher utility for the players. The existence of such equilibria poses an opportunity for players to try and reach cooperative outcomes and increase their payoffs (at the expense of the auctioneer). However, for almost a decade, no natural process has been described, to the best of our knowledge, that attains any of these equilibria. Moreover, in [53], it is shown that even a scenario where players can manipulate the objective functions of their learning algorithms is not sufficient to reach any form of cooperation. In fact, convergence of standard no-regret agents to the second-price outcome induces a dominant strategy for users to provide the true objectives for their agents. To understand how cooperative outcomes could perhaps still be reached, we observe that in the collusive equilibria described in [35], the low-value player wins in some fraction of the auctions and has positive utility, whereas under simple no-regret dynamics (as those studied in [53]), only a single player wins in the limit. This suggests that payments between the agents during their dynamics could enable the agents to share the welfare and reach cooperative outcomes.
@@
The single-item first-price auction two players with values $v_{1}\geq v_{2}$ and any mean-based bidding agents is not stable, and every equilibrium of the payment game is a strong Pareto improvement (i.e., both players are better off) and has lower revenue for the auctioneer compared to the game without payments.
We provide here an outline of the proof, breaking it into a sequence of lemmas. The proofs of the lemmas are deferred to Appendix D. The first part of the proof is the construction of a unilateral payment policy for the high-value player and the analysis of the resulting dynamics. In this policy, the agent with the higher value pays $\eta>0$ to the low-value agent whenever the latter bids zero. We derive the (unique) mixed Nash equilibrium of the game with this payment policy, with $\eta$ as a parameter (Lemma 1). We then evaluate the utilities of the players in that equilibrium and show that the optimal value of $\eta$ for the high-value player is half the value of the second player (Lemma 2), and that in this case, the players have higher utilities than in the game without payments. We next prove that in the game between the agents, while the agents do not play a mixed equilibrium, and their strategies are correlated, nevertheless, the marginal distributions and payoffs are the same as in the mixed Nash equilibrium we derived (Lemma 3). Finally, using the fact that these utilities for the players are obtained with a unilateral payment by the high-value player, we show that these utilities bound from below the utilities in any equilibrium of the payment-policy game. We now state the lemmas carefully.
@@
If the agents reach the Nash equilibrium utilities, the optimal value of $\eta$ for player $1$ is $v_{2}/2$. With this value of $\eta$, the utilities for the players are $v_{1}-v_{2}+\frac{v_{2}^{2}}{4v_{1}}$ for player $1$ and $\frac{v_{2}}{2}$ for player $2$. The winning frequency of player $2$ is in the interval $[0,3/8]$, with the lower bound being attained as $v_{2}/v_{1}\rightarrow 0$ and the upper bound being attained as $v_{2}/v_{1}\rightarrow 1$.
The result above is for Nash equilibrium. While in principle, no-regret agents may reach an arbitrary CCE and not necessarily this outcome, the next lemma shows that the family of Nash-equilibrium distributions parameterized by $\eta$ captures well the marginal distributions and utilities of the broad family of mean-based [19] no-regret agents in these games.
@@
Consider the game where player $1$’s agent pays $\eta$ to player $2$’s agent whenever the latter agent bids zero. Fix a CCE with marginal distributions $F_{1}$ and $G_{2}$ to which the dynamics of mean-based agents converge with positive probability. Then, $F_{1}(b)=F_{1}^{NE}(b)=\frac{\eta}{v_{2}-b}$, $G_{2}(b)=G_{2}^{NE}(b)=\frac{v_{1}-v_{2}+\eta}{v_{1}-b}$, and the players’ utilities are the same as in the Nash equilibrium.
Remark.
Theorem 2 can be extended to some extent to $n$ players by a reduction to a game between the two top-valued players. In this reduction, all the lemmas extend with essentially no change in their proofs, so the result that first-price auctions are not stable extends as well. See Appendix D.2.1.
@@
Figure 2: Dynamics in first-price auctions.
Figure 2 shows simulation results with Hedge (a version of multiplicative weights [82]) agents in a sequence of $100$,$000$ auctions. The left panel compares the Nash equilibrium theoretical prediction from Lemma 3 with the empirical CDF from the agents’ dynamics with $v_{1}=1,v_{2}=0.5$. The right panel compares the theoretical winning frequencies of the players as specified in Lemma 2 with the empirical frequencies as a function of the ratio of values $v_{2}/v_{1}$ in multiple simulations. The long-term marginal bid distributions and winning frequencies are clearly consistent with the theoretical predictions. We observe similar results with other algorithm variants like follow the perturbed leader and linear multiplicative weights [3]. There is very little variation between simulation instances.
Interestingly, it turns out that these dynamics of the agents with payments recover, in a special case, a known collusive CCE distribution of the auction without payments.
@@
For symmetric bidders with $v_{1}=v_{2}=1$ and $\eta=1/e$, the Nash equilibrium bid distribution of the game with payment $\eta$ is the same as that of the minimum-revenue coarse correlated equilibrium of the standard first-price auction (without payments) given in [35].
@@
Notice, however, that in our game, the players’ utilities are different from those of the CCE without payments. In particular, the payment $\eta$ that results in the this CCE distribution is not optimal for player $1$, so it is not an equilibrium of the payment-policy game.
@@
We are now ready to conclude the proof of Theorem 2 by using the lemmas above to bound from below the utility for the high player in any Nash equilibrium of the payment game. The idea is to use the fact that the game analyzed in our proof is induced by a unilateral payment policy by the high-valued player. In any equilibrium, player $1$ can consider the deviation in which he rejects (pays back) all the payments he received and makes this unilateral payment, leading to a utility of $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$. The full proof is in Appendix D.2.
@@
To evaluate the potential loss in welfare in the game considered above, we can analyze the winning frequencies from Lemma 2. The reduction in welfare is equal to the difference $v_{1}-v_{2}$ times the fraction of times in which the low-value player wins. This reduction in welfare is maximized, as can be verified by numerical calculation, when $v_{2}/v_{1}\approx 0.50959...$, at which point the reduction is approximately $8.96\%$. When considering the reduction in revenue, since the support of the bids is between zero and $v_{2}/2$, almost all the payments are strictly less than $v_{2}/2$, so the revenue is reduced to less than half the revenue of the outcome of the game without payments. Thus, the players manage to improve their utilities at the expense of the auctioneer, with a relatively small loss in efficiency.
4 Single-Player Manipulations

### v2

@@ 4 Auctions with Payments between Learning Agents
To gain a deeper understanding of the incentives of users in online platforms to use payments with their learning agents and the potential impact of such payments, we focus on one class of games that has already been widely analyzed: auctions with automated bidders. We consider the simplest model of first- and second-price auctions, where at each time $t\in[T]$ a single identical item is sold. Each player $i$ has a value $v_{i}$ for the item, and we index the players in decreasing order of their values, $v_{1}\geq v_{2}\geq\dots\geq v_{n}$. At round $t$, each agent submits a bid $b_{i}^{t}$ and the auction mechanism determines the identity of the winner and a price $p^{t}$. The payoff for the winning agent from the auction is $v_{i}-p^{t}$; the other agents get payoff zero. In addition to payoffs from the auction, agents get a payoff according to the payments between the agents (as defined in Section 2). Utilities are additive over game rounds. In the first-price auction, the payment is equal to the highest bid; in the second-price auction, the payment is the second-highest bid. In both cases, the highest bidder wins, and we assume tie-breaking according to the index of the agent (as is done, e.g., in [41]), so that if $i<j$ and $i$ and $j$ both have the same top bid, then $i$ is taken to be the winner. All our analyses are similar for arbitrary tie-breaking rules, but breaking ties in favor of the highest-value agent simplifies the presentation.
4.1 Second-price auctions
@@
The second-price auction has been widely studied [101] (see Section 3 for further references). While the auction is known to be incentive-compatible (or “truthful”), it is also known that truthful bidding forms only one out of many Nash equilibria of the stage game. In the context of learning dynamics, it has been shown that regret-minimizing agents in this auction do not converge to the truthful equilibrium; instead, they reach some degree of tacit collusion [63] with lower revenue to the auctioneer than that obtained under truthful bids.
@@
Specifically, simplifying slightly, second-price auctions have two types of non-truthful equilibria. The first type is “overbidding equilibria,” where the high-value player bids anything above her value and the low-value player bids anything below the high value. The second type is “low-revenue equilibria,” where the high-value player bids anything above the low value and the low-value player bids anything below the low value. [^8] Both types of equilibria arise for any number of players $n$; we describe the two-player case for ease of exposition. The latter type is especially interesting for our setting since any mixture of such equilibria is consistent with regret minimization, but yields high welfare to the players. This suggests that when learning agents exchange payments, the high-valued agent may be able to use payments to influence the dynamics with the other agents to induce a better equilibrium for himself.
@@
We show that, indeed, when agents “trade outside of the mechanism” in this way during their learning dynamics, this kind of collusion emerges from the agents’ interaction in a strong way, potentially leading to zero revenue for the auctioneer.
@@
The single-item second-price auction, where the values of the top two valued players are $v_{1}>v_{2}$, is not stable for any regret-minimizing agents for the players. Furthermore, the payment-policy game of this auction has an $\epsilon$-equilibrium where the players capture the full welfare, and the auctioneer gets zero revenue.
@@
The idea of the proof is, first, to construct a unilateral payment policy for player $1$, parameterized by a single parameter $\epsilon$ that determines the amount that agent $1$ pays to the other agents, such that this policy induces a dynamic where player $1$’s (i.e., the high player’s) agent bids its true value and all the other learning agents bid zero. We then show that for every other player, using the zero-payment policy is in fact a best response, and that the (net) payoff for the high-valued player can be arbitrarily close to the full welfare, so player $1$ cannot improve her payoff by more than $\epsilon$ for any $\epsilon>0$. (Note that with the second-price payment rule, if the agents were not restricted to using finite payments, the players would be locked into a game without an equilibrium, where each player tries to commit to making the highest bid and incentivize the others to bid zero.)
Proof.
(Theorem 1). Consider a repeated second-price auction with a single item sold in every step, and $n$ players with values $v_{1}>v_{2}\geq\dots\geq v_{n}$. Assume that the total payment that a player makes is bounded by some large constant $M>2nv_{1}$. Consider the following profile of payment policies. Every player $j>1$ uses the policy of always paying zero. Choose $\epsilon$ with $0<\epsilon<(v_{1}-v_{2})/n$. The payment policy for player $1$ is specified by the following conditions. (1) Whenever $b_{1}=v_{1}$, make a payment $p_{1j}$ to every player $j>1$ of $\epsilon/(n-1)$ if $b_{j}=0$ and zero otherwise. (2) If agent $1$ bids $b_{1}\neq v_{1}$, then she pays $p_{1j}=\frac{M}{n-1}$ to every other player $j>1$. Note that the maximum payment according to these conditions is $M$, as required. We continue with the following claim.
Claim 1.
Proof.
Claim 2.
Proof.
We begin by looking at auctions with $n>2$. In every deviation by some agent $j>1$, in order to ensure that a bid $b<v_{1}$ is not dominated for agent $1$, agent $j$ must pay agent $1$ an amount of at least $M-v_{1}$. The utility for player $j$ in this case is bounded from above by $v_{j}-(M-v_{1})+\frac{M}{n-1}<2v_{1}-\frac{n-2}{n-1}M<2v_{1}\big{(}1-\frac{n(n-2)}{n-1}\big{)}<0$. That is, in every deviation by agent $j$ in which $j$ can win and get a positive utility from the auction (by incentivizing agent $1$ to bid below $v_{j}$ with some probability), the payment is too large, so the resulting utility is negative. So player $j$ prefers not to make payments and to get a utility of $\epsilon/(n-1)$.
To see this, notice that a bid for which player $2$ pays agent $1$ an amount less than $M$ is still dominated and therefore not played by agent $1$. Suppose that there is a bid $b$ for which agent $2$ pays agent $1$ an amount of $M$. When agent $1$ loses with the bid $b$, she gets a utility of zero. When agent $1$ wins, she has an expected payoff of $v_{1}-\operatorname{\mathbb{E}}[b_{2}|b>b_{2}]$. Denote by $q$ the probability that $b$ is a winning bid for agent $1$. We have $v_{1}-\epsilon\leq q(v_{1}-\operatorname{\mathbb{E}}[b_{2}|b>b_{2}])\leq qv_{1},$ so we have $q\geq 1-\frac{\epsilon}{v_{1}}.$ Without loss of generality, we set $v_{1}=1$, and so agent $1$ wins at least $1-\epsilon$ fraction of the time.
Now, consider agent $2$. By the argument above, agent $2$’s payoff from using bids that are above zero is at most $v_{2}\cdot\epsilon<\epsilon$. Therefore, using such bids gives agent $2$ positive regret compared to bidding zero (when she bids zero, she gets a payment of $\epsilon$). Thus, agent $2$ almost always bids zero. It follows that player $2$ cannot make payments that increase her utility, so the policy of always paying zero is a best response. ∎
We are now ready to complete the proof. With the policy profile that we have, player $1$ gets a utility of $v_{1}-\epsilon$. By taking a sufficiently small value of $\epsilon$, player $1$ gets a utility per time step that is arbitrarily close to $v_{1}$. Since $v_{1}$ is the highest possible welfare in the game and player $1$ gets $\epsilon$ close to it for all $\epsilon>0$, this is an approximate best response for player $1$, and we have an $\epsilon$-Nash equilibrium. The auctioneer’s revenue in this equilibrium is zero, other than vanishing profits during the initial learning phase of the agents. ∎
4.2 First-price auctions
@@
In contrast to the second-price auction, the first-price auction is known not to be incentive compatible. Intuitively, the direct dependence of the utility on the bid requires players to constantly optimize their responses to the bids of the other players. There has thus been significant interest in learning strategies, outcomes, and dynamics in this setting (see Section 3 for references). Interestingly, despite the significant difference from the truthful second-price auction, it is well known (and not hard to see) that the Nash equilibrium of the first-price auction yields the truthful outcome of the second-price auction. It has been shown that mean-based ([22], see Appendix C for the definition) regret-minimization dynamics [63], as well as best-response dynamics [86] can converge only to this outcome.
@@
In [41], it was shown that the first-price auction also has “collusive” coarse correlated equilibria [^9] A generalization of the notion of correlated equilibrium, equivalent to all players having no external regret; also known as the Hannan set. See Appendix C for references. (CCE) with lower revenue for the auctioneer than the second-price outcome, and higher utility for the players. The existence of such equilibria poses an opportunity for players to try and reach cooperative outcomes and increase their payoffs (at the expense of the auctioneer). However, to the best of our knowledge, no natural process has been described in nearly a decade since that result that attains any of these equilibria.
Moreover, in [63], it is shown that even a scenario where players can manipulate the objective functions of their learning algorithms is not sufficient to reach any form of cooperation. In fact, convergence of standard no-regret agents to the second-price outcome induces a dominant strategy for users to provide the true objectives for their agents. To understand how cooperative outcomes could perhaps still be reached by learners, we observe that in the collusive equilibria described in [41], the low-value player wins in some fraction of the auctions and has positive utility, whereas under simple no-regret dynamics (as those studied in [63]), only a single player wins in the limit. This suggests that payments between the agents during their dynamics could enable the agents to share the welfare and reach cooperative outcomes.
@@
(1) The single-item first-price auction with two players and any mean-based no-regret bidding agents is not stable: players prefer to use payments. (2) Every equilibrium of the payment game is a strong Pareto improvement (i.e., both players are better off) and has lower revenue for the auctioneer compared to the game without payments.
Remark.
The first part of Theorem 2, showing that first-price auctions are not stable, extends to any number of players by a reduction to a game between the two top-valued players. In this reduction, all the lemmas extend with essentially no change in their proofs. See Appendix B for further details.
We break the proof into a sequence of lemmas. For ease of readability, we outline the proof here and defer the full formal details to Subsection 4.2.1 at the end of this section. The first part of the proof is the construction of a unilateral payment policy for the high-value player and the analysis of the resulting dynamics. Under this policy, the agent with the higher value pays $\eta>0$ to the lower-value agent whenever the latter bids zero. We derive the (unique) mixed Nash equilibrium of the game with this payment policy, with $\eta$ as a parameter (Lemma 1). We then evaluate the utilities of the players in that equilibrium and show that the optimal value of $\eta$ for the high-value player is half the value of the second player (Lemma 2), and that in this case, the players have higher utilities than in the game without payments. We next prove that in the game between the agents, while the agents do not play a mixed equilibrium and their strategies are correlated, the marginal distributions and payoffs are nevertheless the same as in the mixed Nash equilibrium we derived (Lemma 3). Finally, using the fact that these utilities for the players are achieved through a unilateral payment by the high-value player, we show that they provide a lower bound on the utilities in any equilibrium of the payment-policy game. We now state the lemmas formally.
@@
If the agents reach the Nash equilibrium outcomes, the optimal value of $\eta$ for player $1$ is $v_{2}/2$. With this value of $\eta$, the utilities for the players are $v_{1}-v_{2}+\frac{v_{2}^{2}}{4v_{1}}$ for player $1$ and $\frac{v_{2}}{2}$ for player $2$. The winning frequency of player $2$ is in the interval $[0,3/8]$, with the lower bound attained as $v_{2}/v_{1}\rightarrow 0$ and the upper bound attained as $v_{2}/v_{1}\rightarrow 1$.
The result above is for Nash equilibrium. While, in principle, no-regret agents may play an arbitrary CCE and not necessarily this outcome, the next lemma shows that the family of Nash-equilibrium distributions parameterized by $\eta$ captures well the marginal distributions and utilities of the broad family of mean-based [22] no-regret agents in these games.
@@
Consider the game where player $1$’s agent pays $\eta$ to player $2$’s agent whenever the latter agent bids zero. Fix a CCE with marginal distributions $F_{1}$ and $G_{2}$ to which the dynamics of mean-based agents converge with a positive probability. Then, $F_{1}(b)=F_{1}^{NE}(b)=\frac{\eta}{v_{2}-b}$, $G_{2}(b)=G_{2}^{NE}(b)=\frac{v_{1}-v_{2}+\eta}{v_{1}-b}$, and the players’ utilities are the same as in the Nash equilibrium.
@@
Figure 2: Dynamics of Hedge algorithms with payments between agents in first-price auctions.
Figure 2 shows simulation results with Hedge agents (a version of multiplicative weights [97]) in a sequence of $100$,$000$ first-price auctions with payments between the agents. The left panel compares the Nash equilibrium theoretical prediction from Lemma 3 with the empirical CDF from the agents’ dynamics with $v_{1}=1,v_{2}=0.5$, and agent $1$ pays agent $2$ $\eta=v_{2}/2$ when the latter bids zero, as in Lemma 2. The right panel compares the theoretical winning frequencies of the players as specified in Lemma 2 with their empirical winning frequencies as a function of the ratio of values $v_{2}/v_{1}$ in multiple simulations. The long-term marginal bid distributions and winning frequencies are clearly consistent with the theoretical predictions. We observe similar results with other algorithm variants such as follow the perturbed leader and linear multiplicative weights [5]. There is very little variation between simulation instances.
Interestingly, it turns out that these dynamics of the agents with payments recover, as a special case, the known collusive CCE distribution of the auction without payments from [41].
@@
For symmetric bidders with $v_{1}=v_{2}=1$ and payment of $\eta=1/e$, the Nash equilibrium bid distribution of the game with payment $\eta$ is the same as that of the minimum-revenue coarse correlated equilibrium of the standard first-price auction (without payments) given in [41].
@@
Notice, however, that in our game, the players’ utilities are different from those of the CCE without payments. In particular, the payment $\eta$ that results in this CCE distribution is not optimal for player $1$, so it is not an equilibrium of the payment-policy game.
@@
We are now ready to conclude the proof of Theorem 2 by using the above lemmas to establish a lower bound on the utility of the high-value player in any Nash equilibrium of the payment game. The idea is to use the fact that the game analyzed in our proof is induced by a unilateral payment policy applied by the high-valued player. In any equilibrium, the player $1$ can consider the deviation in which she rejects (pays back) all the payments she received and makes this unilateral payment, which leads to a utility of $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$.
@@
To evaluate the potential loss in welfare in the game considered above (as sometimes a player with a lower value wins), we can analyze the winning frequencies from Lemma 2. The reduction in welfare is equal to the difference $v_{1}-v_{2}$ times the fraction of times in which the low-value player wins. This reduction in welfare is maximized, as can be verified by numerical calculation, when $v_{2}/v_{1}\approx 0.50959...$, at which point the reduction is approximately $8.96\%$. When considering the reduction in revenue, since the support of the bids is between zero and $v_{2}/2$, almost all the payments to the auctioneer are strictly less than $v_{2}/2$, so the revenue is reduced to less than half the revenue of the outcome of the game without payments. Thus, the players manage to improve their utilities at the expense of the auctioneer, with a relatively small loss in efficiency.
4.2.1 Proof of Theorem 2
Proof.
(Lemma 1). We start by assuming that in the game between these agents, parameterized by $\eta$, there is a mixed Nash equilibrium with continuous CDFs, PDFs with full support on $[0,v_{2}-\eta]$, and player $2$ has a positive probability of bidding zero. The validity of this ansatz will be verified shortly.
Consider player $1$’s utility. In equilibrium, she is indifferent between bidding zero and bidding $v_{2}-\eta$, giving us the size of the point mass for player $2$ at zero:
[…]
Player $1$’s utility when bidding $x$ is $u_{1}(x,y)=G_{2}(x)(v_{1}-x)-\chi(y)\eta$, where $\chi(y)$ is an indicator which equals $1$ if player $2$ bids zero (i.e., if $y=0$) and equals zero otherwise. Importantly, it is independent of $x$. Player $1$ is indifferent between all her bids, so $\frac{\partial u_{1}(x,y)}{\partial x}=0$, which gives
$G^{\prime}_{2}(x)(v_{1}-x)-G_{2}(x)=0.$
$c=v_{2}G_{2}(0)=v_{1}-v_{2}+\eta.$
Verifying the consistency of the support (the CDF must be equal to $1$ at the top of the support):
$G_{2}(v_{2}-\eta)=\frac{v_{1}-v_{2}+\eta}{v_{1}-(v_{2}-\eta)}=1.$
Thus, we have a mixed Nash equilibrium with CDFs
$F_{1}(x)=\frac{\eta}{v_{2}-x},\quad G_{2}(x)=\frac{v_{1}-v_{2}+\eta}{v_{1}-x},$
as stated in the lemma.
Proof.
[…]
A straightforward calculation by differentiating this expression with respect to $\eta$ shows that the expectation is maximized when $\eta=v_{2}/2$, in which case its value is $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$, as claimed. With this choice of $\eta$, player 2’s utility is $v_{2}/2$.
The frequency at which the low agent (agent $2$) wins is given by:
[…]
Proof.
Claim 3.
Proof.
Continuing the proof of Lemma 3, suppose, for contradiction, $F_{1}(b^{\prime})>F_{1}^{NE}(b^{\prime})$ for some $b^{\prime}<v_{2}-\eta$. We have that $F_{1}(b^{\prime})>F_{1}^{NE}(b^{\prime})$ implies $u_{2}(b^{\prime})=F_{1}(b^{\prime})(v_{2}-b^{\prime})>F_{1}^{NE}(b^{\prime})(v_{2}-b^{\prime})=\eta=u_{2}(0)$. Therefore, bidding zero is not within the support of player $2$.
In the standard first-price auction (without payments), mean-based agents can converge only to CCEs that include bids above $v_{2}-\eta$ in their support; this follows from [63]. This is a contradiction to our assumption. Therefore, we have $F_{1}=F_{1}^{NE}$.
Next, suppose for contradiction that $G_{2}(b^{\prime})>G_{2}^{NE}(b^{\prime})$ for some $b^{\prime}<v_{2}-\eta$. This implies $u_{1}(b^{\prime})=G_{2}(b^{\prime})(v_{1}-b^{\prime})-\eta G_{2}(0)>G_{2}^{NE}(b^{\prime})(v_{1}-b^{\prime})-\eta G_{2}(0)=v_{1}-v_{2}+\eta-\eta G_{2}(0)=u_{1}(v_{2}-\eta)$. That is, bidding $b^{\prime}$ gives player 1 a higher utility than bidding $v_{2}-\eta$, so the latter is not within the support of player 1, which is a contradiction, since the marginal distribution of player $1$ is $F_{1}$, which is supported on $[0,v_{2}-\eta]$. Thus, overall, we have $F_{1}=F_{1}^{NE}$ and $G_{2}=G_{2}^{NE}$.
Proof.
(Theorem 2.) To prove the theorem, we use the fact that the game analyzed in our lemmas is induced by a unilateral payment policy by the high-value player. In any equilibrium, player $1$ can consider the deviation in which she rejects (pays back) all the payments she received and makes this unilateral payment, leading to a utility of $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$. Therefore, in any equilibrium, player $1$ gets at least this utility, which is higher than the utility of $v_{1}-v_{2}$ that she gets in the game without payments. As for the utility of the low-value player, since the high-value player’s utility is above the second price, player $1$ must use positive payments, otherwise the dynamics of mean-based agents could approach only the second-price outcome (as shown in [63]). Therefore, in the equilibrium of the payment-policy game, the low-value player has positive utility, an improvement over the game without payments.
Having established this. the fact that the revenue decreases in any equilibrium is now straightforward. To see this explicitly, suppose for contradiction that the revenue $R$ is at least the revenue $v_{2}$ of the game without payments. The total welfare (of the players and the auctioneer) is $w=u_{1}+R+u_{2}\leq v_{1}$. We then get $v_{1}+\eta\Big{(}\frac{v_{1}-\eta}{v_{1}}\Big{)}+u_{2}\leq v_{1}$, which is a contradiction, since player $2$ cannot get negative utility in equilibrium.
5 Manipulations by One Player